OpenStudy (anonymous):

What fraction is greater than 2/5?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

3/5 is greater than 2/5... 4/5, 1/1, 6/5, 7/5, 8/5, 9/5, 2/1..... There's lots and lots and lots and lost and lots and lots and lost and lots and lots and lost and lots and lots and lost and lots and lots and lost. Unless im confusing the question?

OpenStudy (anonymous):

what about 1/5?

OpenStudy (anonymous):

No that's less than 2/5 =)

OpenStudy (anonymous):

4/10, 1/10, 5,10?

OpenStudy (anonymous):

4/10 = 2/5 so No. 1/10 = 0.5/5 so No. And is that 5/10? 5/10 is greater than 2/5, Yes.

OpenStudy (anonymous):

thank you thank you <3 :)